OpenStudy (openstudygirl2):

1/5+1/6= help me please i suck at math so i will fan and medal

Awolflover1 (awolflover1):

\[\frac{ 1 }{ 5 }+\frac{ 1 }{ 6}\]

Awolflover1 (awolflover1):

so we have to flip it over

Awolflover1 (awolflover1):

\[\frac{ 5 }{ 1 }+\frac{ 6 }{ 1 }=?\]

OpenStudy (openstudygirl2):

the answer is 11/1?

Awolflover1 (awolflover1):

well you know how to get Least Common Denominator?

OpenStudy (openstudygirl2):

no i suck at math i have no idea?

OpenStudy (openstudygirl2):

???

Awolflover1 (awolflover1):

okay what can go in 5 and 6? 5x6?

Awolflover1 (awolflover1):

30 right?

OpenStudy (openstudygirl2):

uh ya i think

Awolflover1 (awolflover1):

okay so we have to do around way with Least Common Denominator so

OpenStudy (openstudygirl2):

sooo....

Awolflover1 (awolflover1):

\[\frac{ 6}{ 30 }+\frac{ 5 }{ 30 }\]

OpenStudy (openstudygirl2):

10/30?

Awolflover1 (awolflover1):

nope not 10

OpenStudy (openstudygirl2):

oh 11/30 sorry

Awolflover1 (awolflover1):

yep there we go!

OpenStudy (openstudygirl2):

thxxx